# Yuri Balashov

### v1

@@ Abstract
Large Language Models (LLMs) excel in translation among other things, demonstrating competitive performance for many language pairs in zero- and few-shot settings. But unlike dedicated neural machine translation models, LLMs are not trained on any translation-related objective. What explains their remarkable translation abilities? Are these abilities grounded in “incidental bilingualism” (Briakou et al., 2023) in training data? Does instruction tuning contribute to it? Are LLMs capable of aligning and leveraging semantically identical or similar monolingual contents from different corners of the internet that are unlikely to fit in a single context window? I offer some reflections on this topic, informed by recent studies and growing user experience. My working hypothesis is that LLMs’ translation abilities originate in two different types of pre-training data that may be internalized by the models in different ways. I discuss the prospects for testing the “duality” hypothesis empirically and its implications for reconceptualizing translation, human and machine, in the age of deep learning.
@@ 1 Introduction
Large Language Models’ (LLMs) relation to translation is curious, both historically and conceptually. Historically, all the landmark achievements leading up to generative AI were made in the context of machine translation (MT), in the space of three years. This includes the original sequence-to-sequence model (Sutskever et al., 2014), the classical attention mechanism (Bahdanau et al., 2014; Luong et al., 2015) and the transformer (Vaswani et al., 2017), as well as the attendant subword segmentation algorithms (Sennrich et al., 2016; Kudo and Richardson, 2018). Introduced initially to improve the state of the art in MT by 2 or 3 BLEU points, [^2] BiLingual Evaluation Understudy (Papineni et al., 2001)—a popular metric for automatic evaluation of MT quality based on an averaged $n$-gram (usually up to $n=4$) overlap between an MT output and a reference translation. these achievements were immediately adopted in virtually every other area of deep learning, from text to image to video generation and more. Dwarfed and sidelined by the recent AI sensations, translation is rarely mentioned as the original driving force behind them. This is unfair. [^3] Future scholars will hopefully sort it out and restore historical justice. As early as 2020 a leading expert described the flow of intellectual resources through MT as follows: “It is only a slightly exaggerated characterization to say that the recent wave of young and excitable deep-learning researchers moved in, showed improvements through their methods, declared success, and moved to bigger and better things” (Koehn, 2020, 29).
@@
From a more theoretical perspective, it is not so clear why LLMs translate as well as they do, showing competitive results on standard benchmarks for many language pairs in zero- and few-shot settings. Unlike dedicated neural MT models, LLMs are not trained on any translation-related objective. Indeed, they are pre-trained almost exclusively on English—basically the entire Anglocentric Internet—with only a small percentage of non-English content (Blevins and Zettlemoyer, 2022; Lin et al., 2022). Translation is, therefore, an emergent ability of LLMs alongside reasoning, text summarization and generation, and so forth (Wei et al., 2022a). What explains its origin?
@@
The sheer size of LLMs (trillions of parameters) and of their pre-training data (trillions of tokens), as well as the presence of some multilingual content in it (on the order of 3–10% in most cases) are important factors. In comparison, a strong transformer-based neural machine translation (NMT) model requires anywhere from 100M to 10B parameters and a comparable number of aligned sentence pairs in two languages to train to perfection. A three-orders-of-magnitude difference in the model and data size, along with an admixture of multilingual content, suggests that LLMs could probably internalize everything anyone has ever translated, as long as it ended up on the internet in one form or another.
@@
But this is just a starting point. While multilingual pre-training data is key to explaining the translation abilities of LLMs, that data is very heterogeneous in form and content. LLMs may have been exposed to generally available parallel corpora (such as the OPUS collection, WMT benchmarks, United Nations proceedings, and other open repositories of multilingual texts often used to train dedicated NMT engines) during pre-training. They may also have seen foreign language teaching texts, bilingual abstracts of scientific papers, and other trace amounts of “incidental bilingualism”—“needles in a haystack,” as Briakou et al. (2023) put it. And in the last two years they may have learned more from translation queries made to the free versions of the largest models.
@@
It is hard to resist the idea that contemporary LLMs must somehow be able to locate and leverage semantically related monolingual content from the noisy multilingual soup of their training data, and that this ability contributes to their translation performance and other multilingual abilities such as responding to questions or instructions in different languages (see, in particular, Chua et al., 2023; Gao et al., 2025). If that is indeed the case, then the translation abilities of LLMs may have two distinct sources that are learned in somewhat different ways. For the sake of discussion, I will refer to these as Local and Global. “Local learning” makes use of bilingual signals present within a single context window (e.g. an English sentence followed soon by its French translation in the training data). “Global learning,” in contrast, capitalizes on mining semantically related monolingual contents that are spread out over the LLM’s pre-training data.
@@
‘Local’ and ‘Global’ are rough-and-ready labels, and I offer them primarily as a heuristic framework for exploration of potentially fruitful conceptual connections. Due to differences in their architecture, training data, and training objectives, dedicated NMT systems and LLMs perform translation tasks in significantly different ways (see Section 3). If, in addition, LLMs learn to translate in two different ways, Local and Global, then neural-network-based translation in a broader sense (encompassing both traditional NMT models and LLMs prompted for translation tasks) cannot be conceptualized as a unified process. Perhaps this is not surprising given the checkered history of MT and the variety of approaches to translation developed along the way (Section 2), each stage introducing a new understanding of what it means, and takes, to translate. What is different nowadays is the stunning translation quality brought by neural networks and deep learning, which makes the need for understanding their inner workings more pressing.
@@
Notably, human translation is far from being a univocal concept either. This is obvious from a cursory glance at the landscape of modern translation studies (Snell-Hornby, 2006; Venuti, 2013; Munday et al., 2022) highlighting the complexity and subjectivity of human translation and its dependence on various linguistic, cultural, cognitive, environmental, and ergonomic factors. Recent progress in computer-assisted human translation technologies has added even more complexity to the picture (Rothwell et al., 2023; Moorkens et al., 2025). At the end of the day, we humans may translate in more than one way, especially when we are equipped with CAT (computer-assisted translation) tools which extend our cognitive abilities in non-trivial manners (Balashov, 2020), and artificial neural networks may do it differently from us, and even from each other. The overall lesson from the recent developments in translation technologies, both human and machine, is that there may be no such thing as Translation with a capital ‘T’.
@@
The primary goal of this paper, however, is to reflect on the origins of the remarkable translation abilities of LLMs in light of recent research and growing user experience. The plan is as follows. Section 2 provides a brief overview of the history of translation technologies with emphasis on the latest developments. Section 3 compares standard state-of-the-art NMT models with LLMs. (Readers familiar with the material of Sections 2 and/or 3 may skip ahead.) Section 4 reviews recent applications of LLMs to translation and translation-related tasks. In Section 5, I discuss important theoretical work exploring how LLMs perform these tasks. In Section 6, I turn to the question of why LLMs are so good at them and develop a “dualistic” proposal about the origin of LLMs’ translation abilities in two types of pre-training data. Section 7 considers prospects for operationalizing and testing this proposed duality empirically. Section 8 draws broader implications for our rapidly evolving concept of translation in the deep learning era. Section 9 offers concluding remarks.
2 Brief History of Translation Technologies
@@
The development of translation technologies has a rich history, with multiple paradigm shifts over the past decades (Figure 1). Each stage in this history introduced new techniques as well as new conceptualizations of what translation entails. This section provides a brief timeline of major developments from early machine translation efforts up to the current integration of MT and LLMs with CAT tools.
Figure 1: Brief history of translation technologies. Period classification for CAT/TM tools adapted from Sin-wai, 2015, Ch. 1.
2.1 Early Machine Translation: Rule-Based Systems
The idea of using machines to automate translation dates back at least to the 17th century, but modern efforts began in earnest after World War II (Hutchins, 1986). A pivotal moment was Warren Weaver’s 1949 memorandum proposing that computers could be used to translate text via logical and cryptographic methods (Weaver, 1949). Early systems in the 1950s, such as IBM’s demonstration with Georgetown University in 1954 (translating 60 Russian sentences into English), used rule-based approaches heavily relying on customized bilingual dictionaries and complicated sets of morphosyntactic rules in order to transfer language-specific grammatical structure from a source to a target language while attempting, in some cases, to encode the meaning of the source sentence in some sort of interlingual representation.
Rule-based machine translation (RBMT) systems achieved limited success but were labor-intensive to build and brittle outside narrow domains. In 1966, the ALPAC report (ALPAC, 1966) commissioned by major funding agencies in the United States famously concluded that MT had failed to meet the inflated expectations. As a result, funding was reduced, slowing progress for a time. Still, RBMT development continued into the 1970s and 1980s, especially in Europe, Canada, and the Soviet Union, gradually improving the linguistic sophistication of rules.
Human translators, meanwhile, began using early computer-assisted translation (CAT) tools (Sin-wai, 2015), beginning with the first translation memory (TM) software Trados (founded ca. 1984). TM systems allowed human translators to reuse past translations stored in a database: given a new source sentence, the tool retrieves similar previously translated sentences for the translator to adapt. This introduced a new way of thinking about translation: less as composing a new target sentence from scratch, more as retrieving and editing known equivalents.
2.2 Example-Based and Statistical Machine Translation
By the late 1980s and especially the 1990s, the increasing availability of digital bilingual corpora spurred a data-driven revolution in MT. The first stage in this process was the development of example-based machine translation (EBMT) methods (Nagao, 1984). EMBT systems use a database of previously translated examples, where new inputs are matched against this database to retrieve and analogically recombine relevant examples to generate an appropriate translation (Somers, 1999). In concept, EBMT and CAT/TM tools have much in common, and their joint historical development is not accidental (Figure 1, adapted from Balashov, 2025). Both are based on classical string matching and retrieval algorithms implementing substitution, recombination, and interpolation of previous translations. But there is an important difference: CAT tools are designed to aid the human translator who is ultimately responsible for assembling computer-retrieved translation segments into a coherent whole, whereas the EBMT systems were expected to do it on their own—a computationally difficult problem that was never solved except in highly restricted domains (Carl and Way, 2003; Wong Tak-Ming and Webster, 2015). EBMT was eventually abandoned in favor of a different paradigm—statistical machine translation (SMT).
Pioneering work by researchers at IBM (Brown et al., 1988, 1993) reconceptualized translation in terms of probabilistic models: the system “learns” from a large corpus of aligned sentences how likely a particular target sentence is, given a source sentence. In practice, SMT systems broke sentences into phrases and learned translation probabilities for phrase pairs, along with a language model for fluent output. Translation came to be seen as a Bayesian decision process: the best translation is the one that maximizes a probability score based on training data. This marked a significant conceptual departure: translation quality became tied to data coverage and frequency statistics rather than linguistic completeness of rules.
SMT proved very effective. By the mid-2000s, it became the dominant approach, with public systems like Google Translate (launched 2006) using SMT under the hood. Over the years, SMT architectures grew in sophistication (from word-based models to phrase-based and syntax-augmented models), but they had clear limitations, often producing disfluent, incorrect, or fragmented output when the required phrase translations were not well attested in the training data. Nevertheless, SMT brought MT quality to a level useful for many applications, and it reinforced a view of translation as largely “analogical”: piecing together translations from observed examples.
Meanwhile, CAT/TM tools matured and gained widespread adoption in professional translation by the 2000s (Figure 1). Users of CAT tools would see suggestions from their TM databases and also, increasingly, suggestions from MT engines. By the 2010s, the synergy between human translators and computers had led to typical workflows where an MT system pre-translates a text and a human expert post-edits it for accuracy and style. The translation process became a human-machine collaboration, blurring the line between automated and manual translation. Translators began to conceive of their tasks as managing an interplay between their own knowledge and machine outputs—another shift in the practice and the concept of translation (Balashov, 2020).
2.3 Neural Machine Translation
The field of MT underwent another drastic change in 2014–2015 when neural network and deep learning algorithms were successfully applied to translation. In 2014, researchers introduced the first end-to-end neural model for MT (Sutskever et al., 2014) based on an encoder-decoder recurrent neural network that reads a source sentence and directly outputs a translation, with all components trained jointly to maximize the likelihood on a parallel corpus. Soon after, the incorporation of attention mechanisms (Bahdanau et al., 2014; Luong et al., 2015) allowed the model to focus on specific source words while generating each target word, greatly improving accuracy and handling of longer sentences. NMT immediately outperformed SMT on many benchmarks, producing more fluent and contextually appropriate translations. By 2016, industry leaders such as Google and Microsoft had replaced their SMT systems with neural models, marking the official beginning of the NMT era.
Crucially, neural networks brought a new perspective: translation could be achieved by a continuous vector space model that learns to implicitly represent meaning. Unlike SMT’s discrete phrase tables, an NMT model maps sentences into high-dimensional latent representations. Many saw this as a revival of the old idea of interlingua (Vauquois, 1968; Hutchins and Somers, 1992), except that now the “interlingua” is a vector in a neural network, learned automatically. Empirically, multilingual NMT systems (one model translating multiple languages) demonstrated that the network can represent similar sentences in different languages close together in this latent space, enabling zero-shot translation between language pairs never seen explicitly in training (Johnson et al., 2017). This suggested that the neural model was capturing an abstract, language-independent representation of the source text—a fundamentally new way to conceptualize translation as grounded in a shared distributional semantics across languages.
With the introduction of the transformer architecture (Vaswani et al., 2017), NMT received a further boost. The transformer, with its multi-head self-attention layers, allowed training of much deeper and larger networks, scaling up MT models dramatically. State-of-the-art NMT today typically uses transformer models with hundreds of millions of parameters, often trained on a comparable number of sentence pairs. The translation quality from such models on high-resource language pairs (such as English-French) is by many accounts on par with average human translators for factual texts, though they can still make errors, especially on complex or low-resource language inputs.
2.4 Current Trends: Hybrid Approaches and Integration with LLMs
@@
Today’s translation technology landscape is one of convergence and integration. The boundary between human and machine translation has continued to blur. Professional translators frequently use CAT tools integrated with a custom NMT system or a generic service such as DeepL or Google. As already noted, the translator’s role is often to post-edit MT output and curate translation memories, a process that loops human expertise back into improving machine suggestions. This symbiosis has led to significant productivity gains and further reconceptualization of translation as a joint human-AI effort, where creativity and critical judgment come from humans, and speed and consistency from machines.
@@
Most recently, the emergence of LLMs has begun to influence translation workflows. Although NMT systems are still the go-to resource for high-volume, high-speed translation needs (Slator, 2023), LLMs have demonstrated impressive translation capabilities even without being specialized for the task. There are early signs of LLMs being integrated into translation pipelines; for instance, using an LLM to refine or evaluate MT outputs, or to handle difficult cases that NMT systems mistranslate. Conversely, practitioners have explored using NMT engines to augment LLMs, for example, by translating non-English user queries into English before feeding them to an English-centric LLM, and then translating the output back to the original language using a domain-specific MT system. [^4] The rapidly growing literature on using LLMs in translation tasks reveals that, with strategic prompting, LLMs can perform progressively sophisticated operations including, but not limited to: evaluating the quality of translation output, including their own (Kocmi and Federmann, 2023; Lu et al., 2024b); spotting and categorizing translation errors and suggesting corrections (Berger et al., 2024; Feng et al., 2024); automatic post-editing of raw MT output (Raunak et al., 2023; Ki and Carpuat, 2024; Alves et al., 2024); adapting translation output to the specific terminology (Ghazvininejad et al., 2023; Rios, 2024), to a given domain (e.g. pharmaceuticals, oil and gas, IT) (Sia and Duh, 2023; Zheng et al., 2024), and to existing translation memories and other project-, client- or domain-specific instructions and reference materials, often outperforming in these respects more traditional approaches earlier implemented in NMT systems (Moslem et al., 2023; Moslem, 2024; Vieira et al., 2024); generating glossaries of special terms from pairs of source and target documents (Ding et al., 2025); improving the quality of translation in low-resource directions (e.g. Swahili-Japanese) by following a “chain-of-thought” (Wei et al., 2022b) prompt which explicitly requires them to pivot (“Translate this sentence from Swahili to English first; then translate the English output to Japanese,” see Section 4.3); following, with benefit, a human translation workflow (Chen et al., 2024; He et al., 2024), often by engaging LLMs in a prolonged interaction involving pre-translation research, drafting, refining, and proofreading (Briakou et al., 2024). The opportunities in this area are virtually unlimited.
To sum up, translation technologies have evolved from rule-based algorithms through statistical methods to neural networks and now to massive pre-trained language models. Each step changed not only the techniques but also our understanding of what it means, and takes, to translate, from symbolic transfer based on formal linguistics (“hire more linguists!”), to translation as data-driven pattern matching (“Every time I fire a linguist, performance goes up”), to translation as deep representation mapping, and now to translation as an emergent property of general language intelligence.

### v3

@@ Abstract
Large Language Models (LLMs) excel in translation among other things, demonstrating competitive performance for many language pairs in zero- and few-shot settings. But unlike dedicated neural machine translation models, LLMs are not trained on any translation-related objective. What explains their remarkable translation abilities? Are these abilities grounded in “incidental bilingualism” (briakou_searching_2023) in training data? Does instruction tuning contribute to it? Are LLMs capable of aligning and leveraging semantically identical or similar monolingual contents from different corners of the internet that are unlikely to fit in a single context window? I offer some reflections on this topic, informed by recent studies and growing user experience. My working hypothesis is that LLMs’ translation abilities originate in two different types of pre-training data that may be internalized by the models in different ways, Local and Global. “Local learning” makes use of bilingual signals present within a single training context window (e.g. an English sentence followed soon by its Chinese translation in the training data). “Global learning,” in contrast, capitalizes on mining semantically related monolingual contents that are spread out over the LLMs’ pre-training data. The key to explaining the origins of LLMs’ translation capabilities is a continuous iteration between Local and Global learning, which is a natural and helpful consequence of batch training. I discuss the prospects for testing the “duality hypothesis” empirically and its implications for reconceptualizing translation, human and machine, in the age of deep learning.
@@ 1 Introduction
Large Language Models’ (LLMs) relation to translation is curious, both historically and conceptually. Historically, all the landmark achievements leading up to generative AI were made in the context of machine translation (MT), in the space of three years. This includes the original sequence-to-sequence model (sutskever_sequence_2014), the classical attention mechanism (bahdanau_neural_2014; luong_effective_2015) and the transformer (vaswani_attention_2017), as well as the attendant tokenization algorithms (sennrich_neural_2016; kudo_sentencepiece_2018). Introduced initially to improve the state of the art in MT by 2 or 3 BLEU points, [^2] BiLingual Evaluation Understudy (papineni_bleu_2001)—a popular metric for automatic evaluation of MT quality based on an averaged $n$-gram (usually up to $n=4$) overlap between an MT output and a reference translation. these achievements were immediately adopted in virtually every other area of deep learning, from text to image to video generation and more.
@@
From a more theoretical perspective, it is not so clear why LLMs translate as well as they do, showing competitive results on standard benchmarks for many language pairs in zero- and few-shot settings. [^3] For a recent comprehensive overview, see ataman_machine_2025. Unlike dedicated neural MT models, LLMs are not trained on any translation-related objective. Indeed, they are pre-trained almost exclusively on English—basically the entire Anglocentric Internet—with only a small percentage of non-English content (blevins_language_2022; lin_few-shot_2022). Translation is, therefore, an emergent ability of LLMs alongside reasoning, text summarization and generation, and so forth (wei_emergent_2022). What explains its origin?
@@
The sheer size of LLMs (trillions of parameters) and of their pre-training data (hundreds of billions of tokens), as well as the presence of some multilingual content in it (on the order of 3–10% in most cases) are important factors. In comparison, a strong transformer-based neural machine translation (NMT) model requires anywhere from 100M to 10B parameters and a comparable number of aligned sentence pairs in two languages to train to perfection. A three-orders-of-magnitude difference in the model and data size, along with an admixture of multilingual content, suggests that LLMs could probably internalize everything anyone has ever translated, as long as it ended up on the internet in one form or another (morris_how_2025).
@@
But this is just a starting point. While multilingual pre-training data is key to explaining the translation abilities of LLMs, that data is very heterogeneous in form and content. LLMs may have been exposed to generally available parallel corpora (such as the OPUS collection, WMT benchmarks, United Nations proceedings, and other open repositories of multilingual texts often used to train dedicated NMT engines) during pre-training. They may also have seen foreign language teaching texts, bilingual abstracts of scientific papers, and other trace amounts of “incidental bilingualism”—“needles in a haystack,” as briakou_searching_2023 put it. And in the last two years they may have learned more from translation queries made to the free versions of the largest models.
@@
It is hard to resist the idea that contemporary LLMs must somehow be able to locate and leverage semantically related monolingual content from the noisy multilingual soup of their training data, and that this ability contributes to their translation performance and other multilingual abilities such as responding to questions or instructions in different languages (see, in particular, chua_crosslingual_2023; gao_understanding_2025). If that is indeed the case, then the translation abilities of LLMs may have two distinct sources and may learn from these sources in two somewhat different ways. For the sake of discussion, I will refer to these as Local and Global. “Local learning” makes use of bilingual signals present within a single context window (e.g. an English sentence followed soon by its French translation in the training data). “Global learning,” in contrast, capitalizes on making the best use of semantically related monolingual contents that are spread out over the LLM’s pre-training data.
@@
‘Local’ and ‘Global’ are rough-and-ready labels, and I offer them primarily as a heuristic framework for exploration of potentially fruitful connections. Due to differences in their architecture, training data, and training objectives, dedicated NMT systems and LLMs perform translation tasks in significantly different ways (see Section 3). If, in addition, LLMs learn to translate in two different ways, Local and Global, then neural-network-based translation in a broader sense (encompassing both traditional NMT models and LLMs prompted for translation tasks) cannot be conceptualized as a unified process. Perhaps this is not surprising given the checkered history of MT and the variety of approaches to translation developed along the way (Section 2), each stage introducing a new understanding of what it means, and takes, to translate. What is different nowadays is the stunning translation quality brought by neural networks and deep learning, which makes the need for understanding their inner workings more pressing.
@@
Notably, human translation is far from being a univocal concept either. This is obvious from a cursory glance at the landscape of modern translation studies (snell-hornby_turns_2006; venuti_translation_2013; munday_introducing_2022) highlighting the complexity and subjectivity of human translation and its dependence on various linguistic, cultural, cognitive, environmental, and ergonomic factors. Recent progress in computer-assisted human translation technologies has added even more complexity to the picture (rothwell_translation_2023; moorkens_automating_2025). At the end of the day, we humans may translate in more than one way, especially when we are equipped with CAT (computer-assisted translation) tools which extend our cognitive abilities in non-trivial manners (balashov_translators_2020), and artificial neural networks may do it differently from us, and even from each other. The overall lesson from the recent developments in translation technologies, both human and machine, is that there may be no such thing as Translation with a capital ‘T’.
@@
The primary goal of this paper, however, is to reflect on the origins of the remarkable translation abilities of LLMs in light of recent research and growing user experience. The plan is as follows. Section 2 provides a brief overview of the history of translation technologies with emphasis on the latest developments. Section 3 compares standard state-of-the-art NMT models with LLMs. (Readers familiar with the material of Sections 2 and/or 3 may skip ahead.) Section 4 reviews recent applications of LLMs to translation and translation-related tasks. In Section 5, I discuss important theoretical work exploring how LLMs perform these tasks. In Section 6, I turn to the question of why LLMs are so good at them and develop a “dualistic” proposal about the origin of LLMs’ translation abilities in two types of pre-training data and two learning processes. Section 7 considers prospects for operationalizing and testing this proposed duality empirically. Section 8 draws broader implications for our rapidly evolving concept of translation in the deep learning era. Section 9 offers concluding remarks.
2 Translation Technologies: Brief History and Current State
@@
The development of translation technologies has a rich history, with multiple paradigm shifts over the past decades (Figure 1). Each stage in this development introduced new techniques as well as new conceptualizations of what translation entails. For useful historical accounts, the reader is referred to hutchins_machine_1986; jurafsky_speech_2009; sin-wai_routledge_2015; koehn_neural_2020; mercan_evolution_2024; ataman_machine_2025.
Figure 1: Brief history of translation technologies
@@
Today’s translation technology landscape is one of convergence and integration (rothwell_translation_2023; moorkens_automating_2025). The boundary between human and machine translation is rapidly blurring. Professional translators use computer-assisted translation (CAT) tools that are often integrated with a custom NMT system or a generic service such as DeepL or Google. The translator’s role is often to post-edit MT output and curate translation memories, a process that loops human expertise back into improving machine suggestions. This symbiosis has led to significant productivity gains and reconceptualization of translation as a joint human-AI effort (balashov_translators_2020), where creativity and critical judgment come from humans, and speed and consistency from machines.
Most recently, the emergence of LLMs has begun to influence translation workflows. Although NMT systems are still the go-to resource for high-volume, high-speed translation needs (makarenko_list_2025; big_language_solutions_why_2025), LLMs have demonstrated impressive translation capabilities even without being specialized for the task. There are early signs of LLMs being integrated into translation pipelines; for instance, using an LLM to refine or evaluate MT outputs, or to handle difficult cases that NMT systems mistranslate. Conversely, practitioners have explored using NMT engines to augment LLMs, for example, by translating non-English user queries into English before feeding them to an English-centric LLM, and then translating the output back to the original language using a domain-specific MT system.
@@
The rapidly growing literature on using LLMs in translation tasks reveals that, with strategic prompting, LLMs can perform progressively sophisticated operations including, but not limited to: evaluating the quality of translation output, including their own (kocmi_large_2023; fernandes_devil_2023; lu_error_2024); spotting and categorizing translation errors and suggesting corrections (berger-etal-2024-prompting; feng_tear_2024); automatic post-editing of raw MT output (raunak_leveraging_2023; ki_guiding_2024; alves_tower_2024); adapting translation output to the specific terminology (ghazvininejad_dictionary-based_2023; rios_instruction-tuned_2024), to a given domain (e.g. pharmaceuticals, oil and gas, IT) (sia_-context_2023; zheng_fine-tuning_2024), and to existing translation memories and other project-, client- or domain-specific instructions and reference materials, often outperforming in these respects more traditional approaches earlier implemented in NMT systems (moslem_adaptive_2023; moslem_language_2024; vieira_how_2024); generating glossaries of special terms from pairs of source and target documents (ding_enhancing_2025); improving the quality of translation in low-resource directions (e.g. Swahili-Japanese) by following a “chain-of-thought” (wei_chain--thought_2022) prompt which explicitly requires them to pivot (“Translate this sentence from Swahili to English first; then translate the English output to Japanese,” see Section 4.3); following, with benefit, a human translation workflow (chen-etal-2024-iterative; he_exploring_2024), often by engaging LLMs in a prolonged interaction involving pre-translation research, drafting, refining, and proofreading (briakou_translating_2024). The opportunities in this area are virtually unlimited.